OpenStudy (anonymous):

Find the height of a cylinder with radius 2.5 cm and surface area 88.59 cm2. A. 10.0 cm B. 5.0 cm C. 6.28 cm D. 3.14 cm

OpenStudy (anonymous):

if u mean all th surface (lateral + the two bases top and low) so its like this: surface of one base -Sb-; \[Sb=3.14 \times 2.5^{2} = 9.625\] => Sb X 2 = 39.25... 2 Sb - All surface = 49.34 cm2 (this is the lateral surface -Sl-) .............(I) Now we have the Base length -Lb-: \[2.5 \times 2 \times 3.14 = 15.7\] .................(II) LE HIGH IS=> LATERAL SURFACE (I) ON BASE LENGTH (II) : ((H = Sl/Lb = 3.14 cm)) ------------------------------------- the true one is : .................. D. 3.14 cm

OpenStudy (anonymous):

Alright so whats the equation to find this one? @MekkiSimo

OpenStudy (anonymous):